OpenStudy (anonymous):

Rewrite the standard form equation below into slope-intercept form. x – 2y = 3

OpenStudy (anonymous):

what i got was y=5x

OpenStudy (anonymous):

x - 2y = 3 -x -x -2y = -x + 3 divided by -2 on both sides y = (x/2) - (3/2)

OpenStudy (anonymous):

i dont know how to do that...

OpenStudy (anonymous):

I'll give you one: x - 5y = 10 what do you do first

OpenStudy (anonymous):

add 5 to both sides

OpenStudy (anonymous):

No, you want 5 on the left side, and x with some number on the left x is positive on the left side, so subtract both sides by x

OpenStudy (anonymous):

x - x -5y = 10 - x you get -5y = 10 - x

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

x - x = 0, so they cancel

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-5y = 10 -x we want the y to be by itself so what do we do

OpenStudy (anonymous):

divide by -5 both sides

OpenStudy (anonymous):

Yes so we get y = (-x/-5) + (10/-5) that simplifies into what?

OpenStudy (anonymous):

10 - x is the same thing as -x + 10. You want the x to be first

OpenStudy (anonymous):

y= (something) +(-2) i dont understand how to do "(-x/-5)"

OpenStudy (anonymous):

(-x/-5) The negatives cancel, leaving you with (x/5) In the end, you have y = (x/5) -2 (10/-5) = -2

OpenStudy (anonymous):

ok i found the answer on the sheet thankyou

OpenStudy (anonymous):

The answer? I just made this up so you know how to do them yourself :)